Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2008-03-22
ISS On-Orbit Status 03/22/08

STS-123-1J/A Flight Day (FD) 12/13.

Crew sleep/wake cycle today: Sleep 4:00am -12:30pm; wake 12:30pm -4:00am tomorrow. HAPPY EASTER WEEKEND!

After wakeup yesterday (~1:30pm EDT) and before breakfast, CDR Peggy Whitson completed another session with the SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight) experiment software for data logging and filling in questionnaire entries in the experiment's laptop session file on the HRF-1 laptop, as suggested on her discretionary 'job jar' task list. (To monitor the crewmember's sleep/wake patterns and light exposure, Peggy wears a special Actiwatch device which measures the light levels encountered by him as well as his patterns of sleep and activity throughout the Expedition. The log entries are done within 15 minutes of final awakening for seven consecutive days, as part of the crew's discretionary 'job jar' task list.)

FE-2-16 Garrett Reisman performed his third session with the biomed experiment INTEGRATED IMMUNE (Validating Procedures for Monitoring Crew member Immune Function), collecting wet saliva samples first thing in post-sleep. FE-2 Leo Eyharts will have his second IMMUNE saliva collection later today after wake-up (~1:30pm). (IMMUNE protocol requires the collection to occur first thing post-sleep, before eating, drinking and brushing teeth, and all samples are stored at ambient temperature. Along with NUTRITION (Nutritional Status Assessment), INTEGRATED IMMUNE samples & analyzes participant's blood, urine, and saliva before, during and after flight for changes related to functions like bone metabolism, oxidative damage and immune function to develop and validate an immune monitoring strategy consistent with operational flight requirements and constraints. The strategy uses both long and short duration crewmembers as study subjects. The saliva is collected in two forms, dry and liquid. The dry samples are collected at intervals during the collection day using a specialized book that contains filter paper. The liquid saliva collections require that the crewmember soak a piece of cotton inside their mouth and place it in a salivette bag; there are four of the liquid collections during docked operations.)

Right after day's begin, FE-1 Yuri Malenchenko continued supporting the experiment session with the Russian/German TEKh-20 Plasma Crystal-3+ (Plazmennyi-Kristall/PK-3+) payload by activating the turbopump in the Service Module (SM)'s Transfer Compartment (PkhO) for keeping the vacuum chamber (ZB) in the SM Work Compartment (RO) evacuated. Afterwards, Yuri performed the experiment in semiautomatic mode, with some manual control, then closed it down and transferred the collected data from HDD (Hard Disc Drive) for downlink to TsUP-Moscow. The turbopump was deactivated this morning at ~3:55am EDT before sleeptime. (Main objective of PK-3 is to study dust plasma wave propagation and dispersion ratio at a specified power of HF discharge, pressure, and a varied number of particles. The research is performed with particles of 6,81 microns (micrometer) diameter at pressures of 20, 40 & 60 Pa. The main purpose is to study the behavior of the structures when affected by a direct electric field of various amplitudes, and to compare their behavior when affected by low frequency alternating electric field. Yuri's task was to create a plasma-dust cloud with a small void in the center with the help of step-down power decrease of the RF generator. During the experiment the FE-1 performed a total of 3 runs.)

CDR Peggy Whitson and FE-2 Leo Eyharts transferred the ER3 ARIS (EXPRESS-3 Active Rack Isolation System) Rack from the US Lab into the Columbus Module (location COL1A1). (The activities included removal of the CEVIS (Cycle-Ergometer with Vibration Isolation System) from the front of the rack, disconnecting data, power and fluid rack umbilicals, disengaging the ARIS isolation plate from the rack standoff, removing both upper and lower snubber cups and detaching the rack for its transfer to COL1A1, after which the ARIS Rack was connected to COL umbilicals and the rack was successfully checked out.)

Whitson & Eyharts also relocated the MSG (Microgravity Science Glovebox) from the US Lab (loc. LAB1S3) to Columbus (loc. COL1F2) after demating the rack umbilicals and releasing the K-BARs (Knee-Brace Assemblies) and pivot mechanisms. Final checkouts are scheduled for later in the week. (The MSG transfer was originally scheduled for tomorrow (Sunday), but the crew decided to complete the move right away as a get-ahead to gain more time tomorrow.)
Peggy and FE-2-16 Garrett Reisman did an outstanding job of installing & configuring additional flexible ventilation ducting from the Node-2 Stbd CCAA (Common Cabin Air Assembly) diffuser (outlet) to the Node-2 stbd aft return inlet. (In response to slightly higher CO2 concentrations in Node-2, Columbus & Shuttle than in the US Lab (a difference of ~1mmHg), probably as the result of inadequate mixing of ISS air flowing between Lab & Node-2, additional fans on Shuttle & Station were turned on, and the ducting installed by Peggy & Garrett. Whitson is scheduled tonight to conduct a CO2 survey with CDMK (CO2 Monitoring Kit) readings in Shuttle & Station to help the ground determine whether the workaround was successful. The decision was made by the IMMT (ISS Mission Management Team) to keep the 13-ft ducting on the ISS for future options, instead of returning it on 1J/A.)

FE-1 Malenchenko installed the IWIS (Internal Wireless Instrumentation System) accelerometer in the Shuttle for taking structural dynamics data during the upcoming Russian thruster firing (tomorrow, 3/23, at 11:52am-12:05pm EDT). The wireless RSU (Remote Sensing Unit) was temporarily set up in PMA-2 and the IWIS arrangement photo documented. RSU remains On.

The FE-1 conducted the periodic (currently daily) checkout/verification of IP-1 airflow sensors in the various RS (Russian Segment) hatchways, including the FGB-to-Soyuz tunnel, and the FGB-to-Node passageway. (This is especially important when the ventilation/circulation system has to cope with a larger crew on board, currently ten persons, and one of the two Russian SKV air conditioners off (SKV-1).),

In preparation for upcoming urine transfer to the Progress M-63 (28P) Rodnik BV1 water tank, Yuri set up the 'plumbing' gear for the regular six compressor test activations to inflate and pressurize the cargo ship's BV1 tank bladder. (The pressurization of the collapsed bladder of the Rodnik water storage tank is conducted as a leak check, lasting ~4h 30min, preparatory to the liquid waste transfer to the tank for disposal. Each of the two spherical Rodnik tanks consists of a hard shell with a soft membrane (bladder) composed of elastic fluoroplastic. The bladder is used to expel water from the tank by compressed air pumped into the tank volume surrounding the membrane.)

Leo Eyharts readied the equipment for the US PHS (Periodic Health Status) w/Blood Labs assessment, to be performed by him tomorrow for the second time, using the PCBA(Portable Clinical Blood Analyzer). Preparations included an electronic function test and control analysis on the PCBA in preparation for tomorrow's blood analysis activity. (The PHS exam, with PCBA analysis and subsequent clinical evaluation, is guided by special software (IFEP, In-Flight Examination Program) on the MEC (Medical Equipment Computer).

Leo & Garrett, the two rotating ISS-16 Flight Engineers, had another 1h 10min set aside on their schedules for generic handover activities (where crewmembers are scheduled together to complete various designated standard tasks for familiarizing the new station resident with procedures, caveats, etc.).

The crewmembers completed most of their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR), TVIS treadmill (FE-1, FE-2), RED resistive exercise device (CDR, FE-2-16) and VELO bike with bungee cord load trainer (FE-1).

Afterwards, Reisman transferred the crew's exercise data file to the MEC laptop for downlinking, as well as the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

Peggy & Garrett worked in the Airlock on preparations for tonight's EVA-5. The tasks included EMU (Extra Vehicular Mobility Unit) relocation, EMU METOX (Metal Oxide) and LiOH (Lithium Hydroxide) canister replacement, EMU battery replacement, configuring EVA tools, preparing the EL (Equipment Lock), readying the EVA digital camera and charging its battery.

Before going into their sleep period this morning, the joint crew had an hour to review timeline and procedures for EVA-5. Then, EV1 Bob Behnken & EV2 Mike Foreman entered the Airlock and began the 10.2 psi mask prebreathe and overnight campout procedures.

For both crews, workday began today at 12:30pm EDT. Behnken & Foreman are currently on 'Campout' (nachalo desaturatsiy = desaturation start) in the A/L CL (Crewlock), after hatch closure this morning at ~3:30am. The two spacewalkers performed PBA (Portable Breathing Apparatus) mask prebreathe for denitrogenation, while readying their tools & equipment, then depressed the CL from 14.7 to 10.2 psi for their sleep period. After wakeup, ending the 8.5-hr sleep period, the CL hatch was cracked at ~1:25pm for a hygiene break/with mask prebreathe. Around 2:00pm, the hatch was closed again by Peggy Whitson (who supports all Campout preps and post-EVA activities as IV/Intravehicular Crewmember) for EVA preparations in 10.2 psi, followed by EMU purge & prebreathe. Afterwards, Peggy supports CL depressurization and EV1/EV2 egress.

EVA-5 Preview: Nominally, EVA-5 begins at ~5:23pm and lasts approximately 6h 30min, ending at ~11:53pm. Its major objectives are to:

Install KAUs (Keep-alive Umbilicals) on the OBSS (Orbiter Boom Sensor System), including APFR (Articulated Portable Foot Restraint) & stanchion/avionics box;
Stow OBSS on ISS main truss (capturing inboard & outboard striker bars, routing W21 cable harness for KAU heaters, installing OBSS cover, cleaning up);
Install MISSE-6 PEC containers (2) in LWAPA(Light Weight Adapter Plate Assembly) on the Columbus EPF (External Payload Facility), photograph LWAPA (trying using the nominal 1/4' PIP pins, or using 3/16' pins from Node-2 stbd vestibule, or go to contingency tie-down plan to avoid return of MISSE to the Shuttle);
Remove JLP (JEM Experiment Logistics Module Pressurized Section) trunnion covers;
Inspect Stbd SARJ (Solar Alpha Rotary Joint) covers, i.e., divot on cover 18 & photograph covers 16, 6, 8, 9 & 11), and
Clean up worksite & ingress.
Before the spacewalk, Leo & Garrett will set up the SSRMS (Space Station Remote Manipulator System) to support the EVA-5, first for OBSS grapple from the SRMS (Shuttle RMS), then hand off the OBSS to Behnken and Foreman during the KAU installation.

After post-EVA cleanup tomorrow morning, sleeptime for the ISS crew begins at ~4:00am, at ~4:30am for the Shuttle crew.

Transfer Status: Transfer is currently on schedule. As of yesterday (FD11), middeck cargo resupply is 92% complete, return cargo transfer is 69% complete, and overall transfer is 78% complete.

STS-123 Update: The OBSS inspection of the port & starboard WLE (Wing Leading Edge) and the Nose Cap RCC (Reinforced Carbon-Carbon) were completed today. Results will be reported at tomorrow's (FD13) MMT meeting.

ATV Update (Flight Day 13): ATV 'Jules Verne' is at its Parking point, ~2000 km ahead of ISS. Next station-keeping maneuver is scheduled for 3/25 but may be cancelled since the orbit is stable.

ASN-M Update: RSC-Energia successfully switched the ASN-M Satellite Navigation System back to its primary NVM-1 Navigation Computer Module. No further activity planned. ASN-M is critically required for ATV docking, and both NVM computers must be functioning properly.

ISS Crew Sleep Shift Planning: To synchronize the ISS crew's timeline with STS-123/1J/A docking and subsequent docked activities, Peggy's, Yuri's and Leo's wake/sleep cycle has undergone a number of shifts which started on 3/11. For the next three days, the wake/sleep shift schedule is as follows (all times EDT):

FD13 Wake: 12:30pm (3/22) - 4:00am (3/23)
Sleep: 4:00am - 12:30pm (3/23)

Wake: 12:30pm (3/23) - 3:30am (3/24)
Sleep: 3:30am - 12:00pm (3/24)

Wake: 12:00pm (3/24) - 3:30am (3/25)
Sleep: 3:30am - 12:00pm (3/25)

Weekly Science Update (Expedition Sixteen -- Week 22)

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Radiation measurements continue to be performed in the PIRS module. Radiation measurements continue to be performed in the PIRS module. Last memory card exchange took place at 3/14.

ANITA: Completed.

BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.

CFE (Capillary Flow Experiment): Reserve.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): In progress.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): In progress.

EuTEF (European Technology Exposure Facility): The EuTEF platform power feeders were de-activated in preparation of the EVA-3 on 3/18. A thermal clock was triggered and the final power down period was well below it. Status of the various instruments: - DEBIE-2: the instrument showed regular link errors. It was subjected to troubleshooting during last couple of days. No conclusive recovery yet, and currently powered off; - DOSTEL: On-going science acquisition; - EuTEMP: Currently inactive as planned; - EVC: after a successful commissioning, some difficulty was encountered to get nominal images acquisition. Under investigation if the problem comes from EVC itself or from the Columbus HRM (High Rate Multiplexer); - EXPOSE: On-going science acquisition; - FIPEX: FIPEX showed Link Errors and was further troubleshooted this week. But progressively activating the sensors in order to create an overload of the communication channel - tests are on-going; - MEDET: On-going science acquisition; - PLEGPAY: The instrument was successfully re-commissioned on 3/19; - TRIBOLAB: in Stand-by mode, awaiting to start science acquisition after ATV docking on 4/3;

Fluid Science Laboratory (FSL): The FSL Facility is awaiting further troubleshooting after 1J/A departure.

GEOFLOW: Troubleshooting on FSL LAN cable repair was successfully performed. However damage on the MIL Standard bus connector was observed where two pins are missing. The FSL Rack was successfully activated on 3/19 after a swap of the active Columbus MIL Bus A for MIL Bus B. Additional anomalies have been encountered while checking out the ODM (Optical Diagnostic Module). Additional repair activities on the LAN and MIL Bus cables are currently planned for 3/22. Start of GEOFLOW is pending further FSL troubleshooting.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): In progress.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: 'Peggy, thank you for unstowing Garrett's saliva kit and having it ready for him to begin his collections. Garrett, you have nearly completed your early Integrated Immune session. Thanks for your participation in our study. Leo, thanks for your participation in your late Integrated Immune session. You will be performing your final saliva collections and blood draw on the Shuttle ride home. Have a safe trip back and we will see you on the ground.'

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 'Pirs'.

MULTIGEN-1: Completed.

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.


PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): 'Peggy, thanks for completing the additional Actiwatch download. After you complete the third week of requested Sleep logs, you will have two more download activities and an Actiwatch Doff. During your next download activity, you will download and initialize two Actiwatches (CDR and 1J/A) and initialize an Actiwatch for the KARI (South Korean) Space Flight Participant (Yi So-yeon). As always, the additional Sleep Logs are greatly appreciated.'

SOLAR (Solar Monitoring Observatory): Anomaly on platform pointing mode under further investigation. No science acquisition possible so far.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

ULTRASOUND: In progress.

WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): The 4 Experiment Containers (ECs) from Rotor A (0-g condition) were transferred to BIOLAB TCU (Thermal Control Unit) at +4degC on 3/12 to slow down the growth. Their transfer from BIOLAB TCU#1 to the MERLIN cooler unit is currently planned for 3/24. These 4 ECs will be downloaded on 1J/A. The 4 ECs of Rotor B are yet to be retrieved by the crew from the blocked Rotor B, after the 1J/A mission. Anomalies during WAICO-Run#1 imply major impact on science return

CEO (Crew Earth Observation): Ongoing.

No CEO photo targets uplinked for today.

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use